OpenStudy (anonymous):

The geometric mean of x and 5 is 15. find the value of x?

myininaya (myininaya):

In equation form: \[\sqrt{x \cdot 5 } =15\]

myininaya (myininaya):

do you know how to solve that for x?

OpenStudy (anonymous):

would it be 3 or 5 square root 3 ?

OpenStudy (apoorvk):

Geometric mean 'G' of 'n' number \(a_1,a_2,a_3,......,a_n\) is: \[G =\sqrt[n]{a_1\times a_2 \times a_3 \times...... \times a_n}\]

OpenStudy (apoorvk):

Here you have 2 nos., 'x' and '5', whose GM is '15'. So, \(15 = \sqrt{x \times 5}\) Square both sides, 15 x15 = x * 5 Can you find 'x' from this?

OpenStudy (anonymous):

so x is 3?

OpenStudy (apoorvk):

Nope. Divide both sides by '5' to calculate 'x'.

OpenStudy (anonymous):

so x is 5 square root 3?

OpenStudy (apoorvk):

Umm no. \[15 \times 15 = x \times 5\] \[or,\frac {225}{5} = x\] Now can you find 'x'?

OpenStudy (anonymous):

no , i can't !!!!

OpenStudy (anonymous):

jk , it's 45 , right ?

OpenStudy (apoorvk):

Wrong!!!

OpenStudy (apoorvk):

Jk, It's right!! :D